# AIGC for Industrial Time Series: From Deep Generative Models to Large Generative Models

### v1

@@ IV-E Privacy Protection
Beginning with AR models, Palekar et al.[31] present a privacy-preserving authentication model for industrial devices. This model employs the Autoregressive Poor and Rich Optimization (APRO)-based secret key generation and Box-Cox transformation for encryption, achieving secure data exchange with minimal computation time and enhanced encryption quality. The study[32] proposes a privacy-preserving framework for smart power networks, integrating enhanced-proof-of-work blockchain and VAE to safeguard data privacy and detect anomalies, while Almaiah et al.[105] introduces a blockchain-based deep learning framework for industrial, employing VAEs with Bidirectional Long Short-Term Memory networks for robust privacy protection. A KingFisher framework[103] uses VAEs for scrutinizing network interactions and safeguarding privacy. Lastly, the privacy-enhanced federated learning mechanism for the Internet of Medical Things[110], utilizes VAEs with differential privacy noise. As for GAN models, Chen et al.[33] discuss a differentially private GAN model to preserve the privacy of machine operation data. Fed-LSGAN[106] integrates federated learning with the least squares GANs for privacy-aware renewable energy scenario generation. Furthermore, the study[104] proposes a solution for debugging data issues in privacy-sensitive and federated learning contexts using differentially private federated GANs. A deep differential privacy data protection algorithm for industrial networks[108] is constructed. Finally, ProcessGAN[107] innovatively generates confidential synthetic process data, especially beneficial for small medical datasets. Besides, the differentially private DDPM [109] is a novel approach integrating differential privacy into diffusion models to ensure data privacy while generating high-quality synthetic data.
@@ V-A Anomaly Detection
In the industrial field, anomaly detection refers to the identification of irregular or abnormal patterns within industrial data, which could signify issues such as equipment failure, production inefficiencies, or security breaches. It is crucial for maintaining operational efficiency, ensuring safety, and minimizing downtime in industrial settings. Advanced analytical techniques, including machine learning and deep learning, are employed for detecting these anomalies by analyzing sensor data, network traffic, or production metrics. For instance, the AE-NAR model[34] is specifically designed for wind turbine pitch-bearing anomaly detection. It leverages an autoencoder for global feature extraction and nonlinear autoregression with augmented Lagrangian for noise reduction, thereby significantly enhancing detection accuracy. Similarly, the adaptive weighted loss VAE[97] is an unsupervised model presented in industrial production for anomaly detection. This model achieves superior accuracy over traditional VAE by adaptively adjusting loss function weights during training. The conditional GAN[87] with its unique encoder-decoder structure adeptly processes high-frequency time series data for anomaly detection in smart manufacturing. These innovative approaches, along with other GANs and Diffusion models[75][27][74], demonstrate the evolving landscape of anomaly detection in industrial. They emphasize the significance of complex computational models in tackling the intricate challenges of anomaly detection within industrial contexts.
@@ V-B Predictive Maintenance
Predictive maintenance is a strategic approach that employs data analytics, machine learning, and advanced computational models to predict and prevent equipment failures before they occur. This approach aims to optimize maintenance operations by identifying potential issues early, thereby reducing unplanned downtime and extending equipment life. A GAN-based data augmentation method[36] is illustrated to address the scarcity of fault condition data in Industry 4.0. This method enhances the training and performance of machine learning models, which are critical for accurately predicting equipment failures. The study[96] further expands on this by introducing a novel approach using Gaussian process regression and GAN for handling missing data, which is vital for making accurate predictions in quality assurance and maintenance scheduling. Additionally, Xiong et al.[91] present the Controlled Physics-Informed GAN (CPI-GAN), a hybrid framework that synthesizes degradation trajectories for enhancing Remaining Useful Life (RUL) predictions. This approach ensures accurate predictions that adhere to fundamental physics, thereby improving maintenance scheduling. In addition to the aforementioned approaches, a Variational Autoencoders (VAEs) method [28] further diversifies the computational strategies employed in the realm of predictive maintenance
@@ V-C Fault Diagnosis
Fault diagnosis refers to the process of identifying, analyzing, and rectifying faults in industrial equipment and systems. It leverages advanced data analytics, machine learning algorithms, and sensor data to accurately and promptly diagnose equipment failures. Zhao et al.[17] introduces the NCVAE-AFL framework for the bearing-rotor system, addressing class imbalance and long-tail distribution issues in fault diagnosis. A data-augmentation approach using time-varying autoregressive models[90] is constructed specifically for fault diagnosis in non-stationary multivariate time series, demonstrating enhanced performance in real-world scenarios. Additionally, the studies[94][37] explore the use of Wasserstein Generative Adversarial Networks (WGAN) for data enhancement, targeting the improvement of fault diagnosis accuracy by addressing the scarcity of mechanical fault samples. A GAN-GDA model[95], combining GAN with gaussian discriminant analysis for data augmentation, aims at enhancing the effectiveness of fault diagnosis in industrial settings. Together with a federated learning method[73], and an AR-based method[111], these studies exhibit the diversity of advanced computational techniques in fault diagnosis. They underscore the crucial role of machine learning and data analytics in effectively diagnosing faults in the industrial environment.
Figure 4: From DGMs to LGMs, the research methods of generative models has undergone profound changes. Building large generative models from four aspects: Large industrial dataset, LGMs architecture for complex industrial characteristics, self-supervised training for industrial time series, fine-tuning for industrial downstream tasks.[112]
@@ V-D Supply Chain Optimization and Production Scheduling
Supply chain optimization and production scheduling are key concepts in operations management, focused on enhancing the efficiency and effectiveness of the entire supply chain through rational resource allocation, coordination of production activities, and process improvement. Recent advancements in this domain have seen the introduction of innovative deep-learning models, particularly leveraging GANs and VAEs. A Conditional Variational Autoencoder (CVAE)[113] based model has been proposed to integrate production scheduling in the automotive component manufacturing sector. This model generates operation-specific health indicators from time series data, aiding in quantifying machine degradation and enhancing production efficiency. The study[89] aims to optimize supply chain management by leveraging deep generative models and time series analysis. This approach is enhanced through the generation of creative alternatives using GANs, thereby aiming to increase the success rate of product introductions. A hybrid model combining TimeGAN and CNN-LSTM networks[93] addresses short-term load forecasting in commercial buildings, improving forecasting accuracy and aiding energy management in supply chains. Finally, the Seq2Seq-WGAN model[92] employs a sequence-to-sequence-wasserstein GAN approach to better capture temporal dependencies in cement production processes. This model generates extensive f-CaO label data and Seq2Seq for managing unequal length input-output sequences, leading to more accurate predictions and enhancing cement industry production scheduling.
@@ VI Large Generative Models for Industrial Time Series
To address these issues, researchers have turned to large generative models such as ChatGPT[11] and DALL·E 2[10]. These models exhibit superior data generation, zero-sample generalization, and multitasking capabilities. The generative AI research is shifting from the traditional paradigm of specific-domain, specific-task, and specialized generation ability to the new paradigm of multi-domain, multi-task, and generalized generation ability. However, there is still no definitive answer on how to construct the large generative model in industrial time series. To promote the research and application of large generative models in industry, we explores how to build large generative models suitable for industrial applications, as illustrated in Fig. 4.
@@ VI-A Large-Scale Industrial Dataset
Large-scale datasets are the cornerstone to drive large generative model research. Unlike the NLP or the CV, the industrial time series are generally sensor readings of time series data, such as pressure, temperature, speed, vibration signals, and power signals. There are a number of open-source industrial datasets of various sizes and domains, such as bearing failure dataset[114], wind turbine dataset[115], three-phase motor failure dataset[116], turbofan engine degradation dataset[117], gearbox failure dataset[118], etc. Specifically, the CWRU[119] dataset is a classical fault diagnosis dataset that records bearing vibration data under different operating conditions, including parameters such as different rotational speeds, loads, and operating times. It consists of four different failure modes, including inner ring failure, outer ring failure, and rolling element failure. The CMAPSS dataset[120] records the turbofan engine degradation monitoring data, which contains six operating conditions and 2 failure states. Further, the recently open-sourced N-CMAPSS dataset[117] contains millions of aircraft engine run-to-failure degradation trajectories, including 128 units and 7 different failure modes. It can be seen that the industry has rich data resources, which provide a solid foundation for building large generative models of the industrial time series.
Although LGMs have accomplished impressive results in the NLP and CV, industrial data involves various types of multi-source devices and sensors with complex device information. Therefore, data processing methods, feature extraction methods, and self-supervised representation learning methods need to be designed according to the unique characteristics of industrial data in order to make the LGMs applicable in this field. There has been some research on such algorithms, for example, an adaptive sensor weighting method[121] based on time-varying Gaussian encoders has been proposed to construct representative features of industrial time series. The article[122] proposes an LSTM-DeepFM model for solving industrial soft sensor problems, which utilizes an LSTM-autoencoder in the pre-training phase to enhance the feature construction process.
@@ VI-B LGMs Architecture Tailored for Industrial Characteristics
Currently, the most representative LGMs include GPT-4[123] and DALL·E 2[10], of which the linguistic generative model GPT-4 adopts Transformer as its core architecture, while the multimodal generative model DALL·E uses the Diffusion Model. Transformer has powerful sequence modeling capabilities, and its self-attention mechanism allows it to process time series data in parallel while efficiently capturing important feature information. Transformer has also made notable achievements in industrial time series, including applications such as remaining life prediction of aircraft engines[112], fault diagnosis of bearings[124], and prediction of degradation processes in batteries[125]. The basic idea of the Diffusion Model[19] is a forward diffusion process to systematically perturb the data distribution and then recover the data distribution through a reverse diffusion process. It is powerful in handling probability distributions and generating samples and is gradually becoming a new paradigm for data generation in industrial time series.
One of the keys to deal with industrial time series data is efficient modeling of complex systems and extracting time series dependencies. For LGMs Architecture, future research could focus on optimizing attention mechanisms, and decomposing and transforming time series data to improve representation and modeling capabilities. These methods include Fast Fourier Transform (FFT)[126, 127], optimising attention mechanisms [128, 121], wavelet transform[129, 130], etc. The computational efficiency of the model also needs to be considered in the design of large generative architectures to accommodate the requirement for real-time performance and efficiency in industrial environment. GT-MRNet[128] is a lightweight time series reduction transformer that reduces the computational cost and the number of parameters by adaptively eliminating redundant time steps.
@@ VI-C Self-supervised Training of Industrial Time Series
The industrial environment involves a variety of heterogeneous devices that may provide different types of data, including temperature, pressure, flow rate, etc. One challenge of self-supervised learning is to process heterogeneous device information and integrate it into a consistent representation. This enables the model to understand data from different devices and achieve pattern generalization. Therefore, domain adaptation and transfer learning approaches can be employed to process heterogeneous information from multiple sources in order to learn domain-independent generalized feature representations. For example, Ren et al.[131] proposed cross-domain learning methods of multi-source black-box data. It achieved cross-domain common mechanism learning in privacy and security scenarios without touching the source domain task data. The paper[132] proposed a supervised generalization method for industrial time series analysis models to extract information from a variety of industrial scenarios.
@@ VI-D Fine-tuning for Industrial Downstream Tasks
Industrial generative models still need to make efforts in multi-task adaptation, and the main fine-tuning methods for large models are as follows: (1) Task-oriented model fine-tuning[134]: this approach typically builds on a pre-trained model and uses task-specific data for further training. This can be achieved by tuning the model parameters, adding new layers, or modifying the loss function. Through this fine-tuning, the model will learn task-specific fine-grained features and representations. (2) Model fine-tuning with prompt learning[135]: Designing templates that fit the upstream pre-training task taps the potential of the pre-trained model, allowing the model to perform the downstream task better with as little labeled data as possible. (3) Model fine-tuning based on adapter networks[136]: adapter networks are lightweight network structures that can be inserted into pre-trained models for learning task-specific representations. Adapters can add additional task-related information to the model without destroying the knowledge of the pre-trained model.
VII Challenges, Future Directions, and Conclusion
VII-A Generative Models for Low-quality and Mixed-type Data

### v2

@@ IV-E Privacy Protection
Palekar et al. [31] introduced an AR-based authentication model (APRO), combining secret key generation and Box-Cox transformation for secure, efficient encryption in industrial devices. The KingFisher framework [100] utilizes VAEs to analyze network interactions and enhance privacy, while a federated learning mechanism [107] applies VAEs with differential privacy noise for secure medical IoT data. In GAN-based methods, Chen et al. [33] proposed a differentially private GAN to protect machine operation data, while Fed-LSGAN [103] integrates federated learning with LSGANs for privacy-aware energy scenario generation. Recently, a differentially private DDPM [106] has been developed, incorporating differential privacy into diffusion models to generate high-quality synthetic data while preserving privacy.
@@ V-A Anomaly Detection
In industry, anomaly detection identifies irregular patterns in data, signaling equipment failures, inefficiencies, or security risks. It is essential for efficiency, safety, and minimizing downtime, using machine learning and deep learning to analyze sensor data, network traffic, and production metrics. The AE-NAR model enhances wind turbine pitch-bearing anomaly detection by combining an autoencoder for global feature extraction with nonlinear autoregression for noise reduction [34]. The adaptive weighted loss VAE improves accuracy in industrial anomaly detection by dynamically adjusting loss weights [90]. Similarly, conditional GANs process high-frequency time-series data in smart manufacturing [78]. Additionally, Chen et al. propose a two-stage adversarial Transformer-based method, integrating an autoregressive TCN and adversarial training, achieving superior performance on real-world industrial datasets [85]. Alongside other GANs and Diffusion models [67, 27, 63], these approaches highlight the growing role of complex computational models in industrial anomaly detection.
@@ V-B Predictive Maintenance
Predictive maintenance is a strategic approach that employs data analytics, machine learning, and advanced computational models to predict and prevent equipment failures before they occur. This approach aims to optimize maintenance operations by identifying potential issues early, thereby reducing unplanned downtime and extending equipment life. A GAN-based data augmentation method addresses fault data scarcity in Industry 4.0, improving machine learning models for failure prediction [36]. Expanding on this, Gaussian process regression with GAN enhances missing data handling for quality assurance and maintenance scheduling [89]. For RUL prediction, CPI-GAN synthesizes degradation trajectories, integrating physics-informed constraints for accurate maintenance scheduling [82]. Additionally, VAEs contribute to predictive maintenance by diversifying computational strategies [28]. Furthermore, Diffusion models improve predictive maintenance by enabling stable industrial time series generation. SDE diffusion with PC samplers enhances reliability [65], while Diff-MTS integrates Ada-MMD and TDR-UNet for superior temporal pattern learning, outperforming GAN-based methods [66].
@@ V-C Fault Diagnosis
Fault diagnosis refers to the process of identifying, analyzing, and rectifying faults in industrial equipment and systems. It leverages advanced data analytics, machine learning algorithms, and sensor data to accurately and promptly diagnose equipment failures. Zhao et al. propose NCVAE-AFL to address class imbalance in bearing-rotor fault diagnosis [17]. A time-varying autoregressive model enhances fault detection in non-stationary multivariate time series [81]. WGAN-based approaches improve fault diagnosis accuracy by tackling mechanical fault data scarcity [87, 37], while GAN-GDA integrates Gaussian discriminant analysis for industrial fault detection [88]. Furthermore, Diffusion models enhance fault diagnosis by generating unknown condition fault signals (DiffUCD [64]), improving feature diversity (PSDM [68]), and creating realistic fault data for digital twin diagnostics [69]. A diffusion-based augmentation method ensures stable sample generation [93]. These studies emphasize the key role of machine learning and data analytics in industrial fault detection.
@@ V-D Supply Chain Optimization and Production Scheduling
Supply chain optimization and production scheduling are key concepts in operations management, focused on enhancing the efficiency and effectiveness of the entire supply chain through rational resource allocation, coordination of production activities, and process improvement. Recent advances leverage GANs and VAEs for deep-learning-based industrial optimization. CVAE aids automotive manufacturing by generating health indicators for machine degradation and efficiency [108]. Deep generative models optimize supply chain management by generating creative alternatives [80]. A TimeGAN-CNN-LSTM hybrid model improves short-term load forecasting for energy management [84], while Seq2Seq-WGAN enhances cement production scheduling by capturing temporal dependencies and generating f-CaO label data [83].
Figure 4: From DGMs to LGMs, the research methods of generative models have undergone profound changes. The roadmap for building large generative models from four aspects: Large industrial dataset, LGMs architecture for complex industrial characteristics, self-supervised training for industrial time series, fine-tuning for industrial downstream tasks.
@@ VI Large Generative Models for Industrial Time Series
The generative AI research is shifting from the traditional paradigm of specific-domain, specific-task, and specialized generation ability to the new paradigm of multi-domain, multi-task, and generalized generation ability[109, 110], such as ChatGPT[11] and DALL·E 2[10]. However, there is still no definitive answer on how to construct the large generative model in industrial time series. To promote the research and application of large generative models in industry, we explores theoretically how to build large generative models suitable for industrial applications, as illustrated in Fig. 4.
@@ VI-A Large-Scale Industrial Dataset
Large-scale datasets are the cornerstone to drive large generative model research. Unlike the NLP or the CV, the industrial time series are generally sensor readings of time series data, such as pressure, temperature, speed, vibration signals, and power signals. There are a number of open-source industrial datasets of various sizes and domains, such as bearing failure dataset[111], wind turbine dataset[112], three-phase motor failure dataset[113], turbofan engine degradation dataset[114], gearbox failure dataset[115], etc. Specifically, the CWRU[116] dataset is a classical fault diagnosis dataset that records bearing vibration data under different operating conditions, including parameters such as different rotational speeds, loads, and operating times. It consists of four different failure modes, including inner ring failure, outer ring failure, and rolling element failure. The CMAPSS dataset[117] records the turbofan engine degradation monitoring data, which contains six operating conditions and 2 failure states.
Although LGMs have accomplished impressive results in the NLP and CV, industrial data involves various types of multi-source devices and sensors with complex device information. Therefore, self-supervised representation learning methods need to be designed according to the unique characteristics of industrial data in order to make the LGMs applicable in this field. For example, an adaptive sensor weighting method[118] based on time-varying Gaussian encoders has been proposed to construct representative features of industrial time series. The article[119] proposes an LSTM-DeepFM model for solving industrial soft sensor problems, which utilizes an LSTM-autoencoder in the pre-training phase to enhance the feature construction process.
@@ VI-B LGMs Architecture Tailored for Industrial Characteristics
Currently, the most representative LGMs include GPT-4[120] and DALL·E 2[10], of which the linguistic generative model GPT-4 adopts Transformer as its core architecture, while the multimodal generative model DALL·E uses the Diffusion Model. Diffusion model is powerful in handling probability distributions and generating samples. Transformer excels in powerful sequence modeling capabilities, and its self-attention mechanism allows it to process time series data in parallel while efficiently capturing important feature information. Transformer has also made notable achievements in industrial time series[121, 122]. However, one of the keys to deal with industrial time series data is efficient modeling of complex systems and extracting time series dependencies. For LGMs architecture, future research could focus on optimizing attention mechanisms, and decomposing and transforming time series data to improve representation and modeling capabilities.
Handling Dynamic Dependencies with Adaptive Attention Mechanisms: Adaptive attention mechanisms aim to refine how attention is distributed over time-series data, focusing on the most relevant time steps. These mechanisms adjust the attention weights based on the dynamic nature of industrial data, helping capture both temporal dependencies[66] (e.g., recent trends) and spatial dependencies[118] (e.g., multi-sensor correlations) effectively.
Domain-Specific Embeddings: Industrial time-series data requires specialized embedding techniques to better represent features like sensor readings over time. Position embeddings can be adapted to handle irregular time intervals, while patch embeddings[123] can be used to segment time-series data into meaningful subsequences, improving the model’s ability to extract local temporal patterns.
Frequency Domain Transformations: Frequency domain transformations such as Fast Fourier Transform (FFT) and wavelet transforms[124] can be applied to industrial time-series data to highlight periodic patterns or anomalies not easily detectable in the raw time domain. These transformations enable the model to focus on important frequency components, improving predictive accuracy.
@@ VI-C Self-supervised Training of Industrial Time Series
The industrial environment involves a variety of heterogeneous devices that may provide different types of data, including temperature, pressure, flow rate, etc. One challenge of self-supervised learning is to process heterogeneous device information and integrate it into a consistent representation. This enables the model to understand data from different devices and achieve pattern generalization. Therefore, domain adaptation and transfer learning approaches can be employed to process heterogeneous information from multiple sources in order to learn domain-independent generalized feature representations. For example, Ren et al.[125] proposed cross-domain learning methods of multi-source black-box data. It achieved cross-domain common mechanism learning in privacy and security scenarios without touching the source domain task data.
@@ VI-D Fine-tuning for Industrial Downstream Tasks
Industrial generative models still need to make efforts in multi-task adaptation, and the main fine-tuning methods for large models are as follows: (1) Task-oriented model fine-tuning: this approach typically builds on a pre-trained model and uses task-specific data for further training. This can be achieved by tuning the model parameters, adding new layers, or modifying the loss function. Through this fine-tuning, the model will learn task-specific fine-grained features and representations. (2) Model fine-tuning with prompt learning: Designing templates that fit the upstream pre-training task taps the potential of the pre-trained model, allowing the model to perform the downstream task better with as little labeled data as possible. (3) Model fine-tuning based on adapter networks: adapter networks are lightweight network structures that can be inserted into pre-trained models for learning task-specific representations. Adapters can add additional task-related information to the model without destroying the knowledge of the pre-trained model.
VII Evaluation Metrics
The quality of time series generation models is typically evaluated based on four core principles: fidelity, diversity, utility, and computational efficiency. This section systematically reviews relevant evaluation metrics and categorizes their principles and significance.
VII-A Fidelity Evaluation
Fidelity measures the consistency between generated and original time series in terms of statistical properties, dynamic patterns, and local context.
VII-A1 Model-based Methods
Indirectly assess generation quality through downstream models:
  • Discriminative Score (DS)[127]: A 2-layer LSTM classifier is trained to distinguish between real and generated time series. The classification error reflects the realism of the generated time series.
  • Predictive Score (PS)[127]: A time series prediction model (e.g., GRU/LSTM) is trained on generated time series and tested on original data using metrics like MAE, evaluating the preservation of temporal dynamics.
  • Contextual-FID (C-FID)[128]: Extends the FID metric from image generation to time series by computing the Fréchet distance between feature distributions of generated and real time series using context-aware embeddings.
VII-A2 Feature-based Methods
Directly compare statistical feature differences.
  • Marginal Distribution Difference (MDD[129]): Computes the average absolute difference between histograms of original and generated time series, assessing distribution alignment.
  • Correlation Score (CS)[130]: Measures the difference in correlation coefficients between original and generated time series, evaluating the preservation of temporal dependencies.
VII-A3 Distance-based Methods
Provide deterministic metrics to avoid model randomness.
  • Euclidean Distance (ED)[131]: Computes the mean pointwise Euclidean distance between generated and original time series, suitable for normalized data.
  • Dynamic Time Warping (DTW)[66]: Captures similarity in asynchronous temporal patterns by finding the optimal alignment path, robust to temporal shifts.
VII-B Diversity Evaluation
Diversity ensures that generated time series cover the original data distribution and avoid mode collapse:
  • t-SNE/PCA Visualization: Projects high-dimensional time series into 2D space to visually compare the distribution coverage and cluster structures of generated and real data points.
  • Distribution Plot Analysis: Plots kernel density estimates of key statistics (e.g., mean, variance, extremes) to verify the alignment between generated and original data distributions.
VII-C Utility Evaluation
Assesses the practical utility of generated data in real-world tasks:
  • Predictive Performance (PS)[127]: As described in fidelity evaluation, PS also reflects the ability of generated data to support downstream prediction tasks.
  • Downstream Task Validation: We consider adopting the Train on Synthetic and Real, Test on Real (TSRTR) paradigm to train classification/regression models on generated data and evaluate their performance on real test sets, directly validating the practical value of generated data.
VII-D Computational Efficiency Evaluation
  • Training Time: Records the wall-clock time from model initialization to convergence, particularly focusing on scalability for large-scale data.
VII-E Analysis
Model-based methods, such as DS and PS, are the most commonly used for evaluating the fidelity of generated time series. However, these methods are sensitive to model architecture, hyperparameters, and training randomness, which can lead to inconsistent results. To mitigate these issues, multiple runs with result averaging are recommended. While visualization methods like t-SNE/PCA offer valuable insights, they rely on subjective interpretation and may not provide precise quantitative assessments. Comprehensive evaluation should be implemented to provide a more reliable and deterministic evaluation. Ultimately, the gold standard for evaluating industrial time series generation is the effectiveness of the generated data in downstream tasks, assessing its practical performance.
VIII Case Study: Application of DGMs in Engine Predictive Maintenance
In this case study, we implement a type of deep generative model, the diffusion model, to enhance downstream tasks in aircraft engine maintenance. Specifically, we focused on improving predictive models by generating synthetic data and combining it with real-world sensor data.
Figure 5: Workflow of implementing diffusion model in this case study
VIII-A Workflow
As illustrated in Fig. 5, the workflow outlines the steps for deploying a diffusion model-based approach to generate synthetic industrial data for a case study focused on predictive maintenance of aircraft engine sensors.
Step 1: Learning representation of industrial time series in the training phase In the training phase, the process begins with the collection of original engine sensor data, which undergoes initial preprocessing to prepare it for further analysis. This preprocessing step typically involves normalization and cleaning, ensuring that the data is in a suitable format for the model. Once the data is preprocessed, a diffusion process is applied. In this stage, Gaussian noise is progressively added to the original data across multiple time steps, simulating the degradation of the data. This process is mathematically described by:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$
where the noise $\beta_{t}$ is introduced at each time step to corrupt the original data. Following the noise addition, the model is trained to predict and reverse this noise. A denoising model is employed, typically utilizing a U-Net architecture to learn how to recover the original data by progressively removing the noise added during the diffusion process. The model is trained to minimize the difference between the predicted noise and the actual noise added, refining its ability to reconstruct the data accurately.
Step 2: Limited sample augmentation in the inference phase During inference, the trained model starts with random Gaussian noise. Through iterative denoising, the model generates synthetic data that mimics the original engine sensor data. The denoising process is mathematically represented by:
$x_{t-1}=\mu_{\theta}(x_{t},t)+\sigma_{t}z,\quad z\sim\mathcal{N}(0,I)$
where the model predicts the noise at each step and adjusts the data accordingly.
Step 3: Evaluation and industrial application Once the synthetic data is generated, it undergoes quality evaluation to ensure it aligns well with the statistical properties of the original data. Metrics such as fidelity and diversityare used to assess the quality of the synthetic data, where fidelity measures how closely the synthetic data matches the original, and diversity reflects the range of variations within the generated dataset. The synthetic data generated through this process is then used for downstream applications. In this case study, the generated data is specifically applied to aircraft engine predictive maintenance. By using synthetic data, predictive models can be evaluated and improved, ultimately aiding in the early detection of potential engine failures, optimizing maintenance schedules, and improving overall reliability in aircraft engine operations.
VIII-B Experiment Setup
We conducted experiments using the CMAPSS engine datasets, which include four subsets: FD001, FD002, FD003, and FD004. These datasets contain time-series sensor data from aircraft engine systems. We deployed the Diff-MTS model[66] and evaluated its performance against other DGMs and the traditional Synthetic Minority Over-sampling Technique (SMOTE[132]). The evaluation focused on fidelity, diversity, and usability of the synthetic data. Finally, we implemented an industrial downstream task focused on predictive maintenance, evaluating the impact of augmented data on predictive accuracy.
VIII-C Experiment Results
Fidelity of the synthetic data: To evaluate the fidelity of generated time series, we compare Diff-MTS with diffusion-based methods (DiffWave[133]) and GAN-based methods (GAN-LSTM[134]) and SMOTE[132]. Synthetic data is generated with lengths of 48, and fidelity is measured using the discriminative score, where lower values indicate higher similarity to real data. Fig. 6 illustrates the discriminative scores of different methods. Diff-MTS consistently achieves lower scores than GAN-based methods and SMOTE, demonstrating superior fidelity. For the four datasets, Diff-MTS achieves discriminative scores of 0.397, 0.164, 0.386, and 0.159, respectively, outperforming traditional generative methods SMOTE of 0.452, 0.447, 0.468, and 0.468. These results further demonstrate the superior fidelity of Diff-MTS in generating synthetic time series with a closer match to the real data distributions.
Figure 6: Discriminative scores of different generative methods across four datasets. Lower scores indicate higher fidelity.
Diversity of the synthetic data: To assess the diversity of synthetic data, we apply PCA and t-SNE to reduce the dimensionality of real and generated time series data. We compare diffusion-based methods (Diff-MTS) with GAN-based methods and traditional method (SMOTE) by visualizing the distributions in two-dimensional space. Fig. 7 presents the PCA and t-SNE visualizations for different methods. Diff-MTS exhibits broad and well-mixed distributions with real data, indicating high diversity. In contrast, GAN-based methods show noticeable clustering and partial overlap, suggesting limited variability. SMOTE-generated data remain highly concentrated, demonstrating the least diversity. These results confirm that the diffusion-based method generates more diverse synthetic data, effectively capturing the variability of industrial time series.
(a) Diffusion-based method
(b) GAN-based method
(c) Traditional method
Figure 7: The synthetic and original data are visualized using PCA (first column) and t-SNE (second column). Each row displays both visualizations for the three methods. Original samples are depicted in blue, while synthesized samples are shown in brown. Greater overlap between points indicates higher similarity between the two data types.
TABLE III: Performance comparison of predictive maintenance using different data augmentation strategies
Method Data Metrics FD001 FD002 FD003 FD004
mixture
Original RMSE 14.453 22.087 14.488 27.971
dataset MAE 10.617 17.033 10.686 22.305
SMOTE RMSE 23.647 27.481 27.231 34.430
MAE 17.241 22.960 21.851 28.980
Diffusion RMSE 13.688 23.511 13.495 19.421
MAE 9.846 19.539 9.564 15.350
SMOTE RMSE 13.585 19.038 16.683 16.493
MAE 10.149 16.305 12.581 12.815
Diffusion RMSE 13.446 18.341 13.412 16.451
MAE 9.387 15.295 9.921 12.615
Performance of downstream application using limited data augmentation: This experiment leverages an industrial time-series diffusion model (Diff-MTS) to generate synthetic sensor data at three times the volume of the original dataset, which is then combined with the original data to enhance the training set for predictive maintenance tasks. As listed in Table III, comparative methods include the classic oversampling technique SMOTE and a baseline model using only the original data. The evaluation is conducted on aircraft engine sensor datasets (FD001-FD004), with Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as the performance metrics. Here, ”Data mixture” denotes the training strategy that combines synthetic and original data, whereas the absence of mixing indicates the use of synthetic data alone.
The results demonstrate that the synthetic data generated by the diffusion model significantly improves model performance. Under the mixed training strategy, the diffusion method achieves optimal results across all four subsets. For FD001, RMSE is 13.446 (a 7.0% improvement over the original data) and MAE is 9.387 (an 11.6% improvement); for FD004, RMSE and MAE are reduced to 16.451 (a 41.2% improvement) and 12.615 (a 43.4% improvement), respectively. Compared to the SMOTE mixed strategy, the diffusion method outperform it consistently across FD001-FD004. Notably, on the FD004 dataset, which represents complex operating conditions, the diffusion method further reduces RMSE by 19.6% compared to SMOTE, highlighting its superior capability in capturing non-stationary temporal patterns. Further analysis reveals that even when using only the synthetic data generated by the diffusion model (without mixing with the original data), the RMSE (13.688) and MAE (9.846) already outperform those of the original dataset (14.453 and 10.617) on the FD001, indicating the high fidelity of the synthetic data.
IX Challenges, Future Directions, and Conclusion
IX-A Generative Models for Low-quality and Mixed-type Data